# Metric theory of inhomogeneous Diophantine approximations with a fixed matrix

### v2

@@ 3 Main results
Before stating our main results, we would like to mention that the case of trivially singular $\Theta^{\top}$ provides us with "the worst possible" approximation properties for almost all the pairs $(\Theta,\boldsymbol{\eta})$: namely, the form $\Theta{\bf q}-\boldsymbol{\eta}||$ is uniformly bounded from below for all ${\bf q}\in\mathbb{Z}^{m}$. We will formulate a more precise statement in Remark 4 in Section 7.1.
@@
In this section we discuss relations between asymptotic approximations of matrix $\Theta^{\top}$ and uniform approximations of pairs $(\Theta,\boldsymbol{\eta})$; that is, we prove statements of a form:
@@ 3.1.1 Dirichlet’s Theorem (due to Jarník)
We start this section with the revision of classical results. Theorem A (which is a reformulation of Theorem 3 A) follows from the argument by Jarník from seminal paper [23] (see also [24, 25]). This Theorem can be thought of as an analog of the classical Dirichlet’s theorem for a fixed $\Theta$.
@@
Then, any pair $(\Theta,\boldsymbol{\eta})$ is $g\circ\frac{1}{\kappa^{m}}$-Dirichlet, that is,
@@
If we assume in addition that $f$ satisfies the condition (6), it implies that
@@ 3.1.2 Davenport-Schmidt’s Theorem
More specifically, it will be shown that this statement holds in restriction on certain affine subspace $\mathcal{A}$. We note that Theorem 3 B corresponds to the case $\mathcal{A}=\mathbb{R}^{n}$. To formulate this result we need a notion of $\Xi$-exceptional affine subspace. The exact definition is a little bit cumbersome and is related to asymptotic directions for the best approximation vectors. It will be given in Section 6.1. In fact, in Section 6.1 we show the the set of all non $\Xi$-exceptional affine subspace is a set of full measure and give an upper bound for the Hausdorff dimension of the set of $\Xi$-exceptional subspaces in Corollary 3.
@@
If we assume in addition that $f$ satisfies the condition (6), it implies that the set
@@ 3.1.3 The set of singular vectors is winning
Theorem 1 shows in particular that the set of non-singular shift vectors $\boldsymbol{\eta}$ has full Lebesgue measure. Assuming the condition (6), we can strengthen this statement, showing that this set of full measure is also winning in the Hyperplane Absolute Winning game. We prove this winning property in a more general form, in restriction on certain affine subspaces, however now the condition on subspaces is different and involves the set of asymptotic directions of the matrix $\Theta^{\top}$. We introduce the asymptotic directions set and briefly discuss its properties in Section 5.3.
@@
Suppose $f$ satisfies the condition (6), and $\Theta^{\top}$ is an $f$-approximable matrix. Let $\Omega$ be the set of all the asymptotic directions for the best approximations of $\Theta^{\top}$.
@@
A more general version of Theorem 2 (without assuming (6)) is formulated in terms of irrationality measure functions in Theorem 2.
@@ 3.2.1 Asymptotic Dirichlet’s Theorem (due to Jarník)
As in Section 3.1, we start with a general result which follows from Jarník’s arguments. This result can be treated as analogous to the classical asymptotic corollary from Dirichlet’s theorem in the inhomogeneous case with a fixed matrix $\Theta$. It is a general version of the statement of Theorem 4 A.
@@
Suppose $\Theta^{\top}$ is a non $f$-Dirichlet matrix. Then, for any $\boldsymbol{\eta}\in\mathbb{R}^{n}$ the pair $(\Theta,\boldsymbol{\eta})$ is $g\left(\frac{1}{\kappa^{m}}T\right)$-approximable:
@@
If we assume in addition that $f$ satisfies the condition (6), it implies that
@@
3.2.2 BAD is null
The next theorem shows: if not all the vectors $\boldsymbol{\eta}$ belong to $\widehat{\bf W}_{n,m}^{\Theta}[g]$, then the set of $g$-badly approximable shifts $\boldsymbol{\eta}$ has zero Lebesgue measure. We note that in a classical special case $g=f_{1}$ of badly approximable pairs this was shown for non-singular [^3] Non-$f_{1}$ singular in our notation $\Theta$ in [31] (Corollary 1.4).
@@ 3.2.2 BAD is null
has Lebesgue measure zero in $\mathbb{R}^{n}$. In particular, if we assume that the condition (6) holds, then the set
@@
of $g$-Badly approximable shift vectors $\boldsymbol{\eta}$ has measure zero, which is equivalent to part B of Theorem 4.
3.2.3 BAD is winning
Despite being of measure zero, the set $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[g\right]$ is known to have full Hausdorff dimension, and moreover, has a stronger winning property. The dimension and winning properties of these sets were widely studied before. It was shown by Kleinbock in [33] that the set $\widehat{{\bf BA}}_{n,m}[f_{1}]$ of badly approximable pairs has full Hausdorff dimension, and by Bugeaud, Harrap, Kristensen and Velani in [6] that the set $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[f_{1}\right]$ has full Hausdorff dimension on certain fractals. It was shown by Tseng [48] that the set $\widehat{{\bf BA}}_{1,1}^{\Theta}\left[f_{1}\right]$ is winning in Schmidt’s $(\alpha,\beta)$-game. It was then shown by Moshchevitin ([40], Theorem 2) that the set $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[g\right]$ is winning, provided that the matrix $\Theta^{\top}$ is $f$-Dirichlet and $g(T)=\frac{1}{f^{-1}\left(\frac{1}{T}\right)}$. In particular, it implies that $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[f_{1}\right]$ is winning. It was shown by Einsiedler and Tseng in [16] (Theorem 1.4) and also independently by Broderick, Fishman and Kleinbock in [8] (Corollary 1.4) that the sets $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[f_{1}\right]$ are winning on certain fractals, and strengthened by Broderick, Fishman and Simmons in [10] (Theorem 1.1) who proved that the aforementioned set is HAW. We use similar techniques to prove our Theorem 4 which generalizes all these results, and proves the winning properties on certain subspaces. We note that a more general result (without the restriction (6)) is formulated in terms of irrationality measure functions, and can be found in Theorem 2.
@@ 3.2.3 BAD is winning
Suppose $f$ satisfies the condition (6), and $\Theta^{\top}\in{\bf D}_{m,n}\left[f\right]$ is an $f$-Dirichlet matrix. Let $\Omega$ be the set of all the asymptotic directions for the best approximations of $\Theta^{\top}$.
@@
Let us notice that the winning properties of the sets $\widehat{{\bf BA}}_{n,m}^{\Theta}\,\left(=\widehat{{\bf BA}}_{n,m}^{\Theta}\left[f_{1}\right]\right)$ on affine subspaces were studied for the case of badly approximable $\Theta$ in [2]: they proved (Theorem 2.4), in the weighted setting, that the aforementioned sets are winning in the classical Schmidt’s $(\alpha,\beta)$-game on any affine subspace $\mathcal{A}$. One can ask if the winning property on arbitrary affine subspaces holds for the general $\Theta$ (and thus, if our restrictions on $\mathcal{A}$ in Theorem 4 can be omitted). The negative answer to this question is given in [15]; namely, they prove the following statement (Theorem 3.1):
@@ 3.2.4 Kurzweil-type theorems
The necessary and sufficient condition (which is also a convergence-divergence condition for a certain series) is only known when $\Theta\in\mathbb{R}$; see [18]. Theorem 3 provides a condition under which a set $\widehat{\bf W}_{n,m}^{\Theta}[g]$ has full Lebesgue measure. In this section we want to show a (non-optimal) condition under which $m\left(\widehat{\bf W}_{n,m}^{\Theta}[g]\right)$.
@@
  • (a)
    $\sum\limits_{\nu=1}^{\infty}\lambda^{\frac{1}{m}}\left(Y_{\nu}^{n}\right)<\infty,$ (13)
@@
  • (b)
    The function $\frac{\lambda(T)}{f(T)}$ monotonically increases. (14)
@@
$\tilde{g}(T)=\tilde{g}_{n,m}(T)=T^{\frac{n}{m}}\cdot\frac{f^{\frac{n}{m}}\left(H(T)\right)}{H(T)}=\frac{\lambda^{\frac{n}{m}}\left(H(T)\right)}{H(T)}.$ (15)
@@
Theorem 6 follows from Theorem 3 via the identifications
@@
Suppose that $\Theta^{\top}$ is an $f$-Dirichlet matrix, $\lambda$ is a decreasing function on $\mathbb{R}_{+}$, and the conditions (6), (8), (9) and (10) are satisfied.
@@
$\sum\limits_{\nu=1}^{\infty}\left(\frac{u(Y_{\nu}^{n})}{g(Y_{\nu}^{n})}\right)^{\frac{\gamma}{m+\gamma n}}=\sum\limits_{\nu=1}^{\infty}\lambda^{\frac{1}{m}}\left(Y_{\nu}^{n}\right)<\infty,$ (16)
@@
Let us note that we can always find such a $u$ that satisfies conditions (8) and (16): one can always take
@@
for $\varepsilon>0$. The convergence in (16) is guaranteed by Proposition 1 A.
@@ 3.3 Khintchine’s results revisited
For any pair $(\Theta,\boldsymbol{\eta})$ that is not trivially singular one has $\omega(\Theta)\geqslant\hat{\omega}(\Theta,\boldsymbol{\eta})$, meanwhile for any $\Theta$ there exists $\boldsymbol{\eta}$ with $\hat{\omega}(\Theta,\boldsymbol{\eta})=\omega(\Theta)$. In another words
@@
In Section 9 we formulate and prove a general result in terms of irrationality measure functions (Theorem 1) which leads to Theorem 8. Here we would like to formulate some comments to Theorem 8.
@@
It is natural to ask whether it is possible to generalize the statement B of Theorem 8 on affine subspaces of $\mathbb{R}^{n}$. Here, we want to show why this is not always possible in a simple low-dimensional case. One can easily construct several straightforward higher dimensional generalizations of this statement.
@@
Indeed: let us show that we can find such an $\mathcal{A}$ of the form $\mathcal{A}_{\eta_{1}}=\{(\eta_{1},\eta):\,\,\eta\in\mathbb{R}\}$, where we can take $\eta_{1}$ from a HAW subspace of $\mathbb{R}$.
@@
is HAW. Fix such an $\eta_{1}$; it is easy to see that, just by the simple observation that
@@ 4.1 An example: powers and logarithms
$f_{\bf a}=f_{(a,b)}=T^{-a}\left(\log T\right)^{-b}.$ (18)
Then, according to (2), the inverse function of $f_{\bf a}$ is equal to $g(T)=(1+o(1))\cdot g_{\bf a}(T)$, where
$g_{\bf a}(T)=a^{-\frac{b}{a}}T^{-\frac{1}{a}}\left(\log T\right)^{\frac{b}{a}}.$ (19)
Let us also notice that due to Proposition 1 A, for any $\delta>0$ one can take $\lambda(T)=\left(\log T\right)^{-m(1+\delta)}$ in Theorem 6, and in this case the function $\tilde{g}$ in Theorem 6 satisfies
$\tilde{g}(T)\asymp\tilde{g}_{{\bf a},\delta}(T)=T^{-\frac{1}{a}}\left(\log T\right)^{\frac{b}{a}-\frac{(m+n)(1+\delta)}{a}}=\left(\log T\right)^{-\frac{(m+n)(1+\delta)}{a}}\cdot g_{\bf a}(T).$ (20)
Note that we can make the related constants as small as we want dividing $\lambda$ by a constant.
@@
In a similar vein, we show some results for the asymptotic inhomogeneous approximation:
@@
4.2 Criteria of badly and very well approximability
In a recent paper [17], the following two criteria for bad approximability and very well approximability were presented:

### v3

@@ 3 Main results
Before stating our main results, we would like to mention that the case of trivially singular $\Theta^{\top}$ provides us with "the worst possible" approximation properties for almost all the pairs $(\Theta,\boldsymbol{\eta})$: namely, the form $\Theta{\bf q}-\boldsymbol{\eta}||$ is uniformly bounded from below for all ${\bf q}\in\mathbb{Z}^{m}$. We will formulate a more precise statement in Remark 5.
@@
In this section we discuss relations between asymptotic approximations of matrix $\Theta^{\top}$ and uniform approximations of pairs $(\Theta,\boldsymbol{\eta})$; that is, we prove statements of the form:
@@ 3.1.1 Dirichlet’s Theorem (due to Jarník)
We start this section with the revision of classical results. Theorem A (which is a reformulation of Theorem 3 (a)) follows from the argument by Jarník from a seminal paper [25] (see also [26, 27]). This Theorem can be thought of as an analog of the classical Dirichlet’s theorem for a fixed $\Theta$.
@@
Then, any pair $(\Theta,\boldsymbol{\eta})$ is $g\circ\frac{1}{\kappa^{m}}$-Dirichlet (where $\kappa$ is defined in (1.3)), that is,
@@
If we assume in addition that $f$ satisfies condition (2.2), then
@@ 3.1.2 Davenport-Schmidt’s Theorem
More specifically, it will be shown that this statement holds in restriction on certain affine subspace $\mathcal{A}$. We note that Theorem 3 (b) corresponds to the case $\mathcal{A}=\mathbb{R}^{n}$. To formulate this result we need a notion of $\Xi$-exceptional affine subspaces (here $\Xi$ is some positive real parameter). The exact definition is a little bit cumbersome and is related to asymptotic directions for the best approximation vectors. It will be given in Section 6.1. In fact, in Section 6.1 we show the the set of all non $\Xi$-exceptional affine subspace is a set of full measure and give an upper bound for the Hausdorff dimension of the set of $\Xi$-exceptional subspaces in Corollary 3.
@@
If we assume in addition that $f$ satisfies condition (2.2), then the set
@@ 3.1.3 The set of singular vectors is winning
Theorem 1 shows in particular that the set of non-singular shift vectors $\boldsymbol{\eta}$ has full Lebesgue measure. Assuming condition (2.2), we can strengthen this statement, showing that this set of full measure is also winning in the Hyperplane Absolute Winning game. We prove this winning property in a more general form, in restriction on certain affine subspaces, however now the condition on subspaces is different and involves the set of asymptotic directions of the matrix $\Theta^{\top}$. We introduce the asymptotic directions set and briefly discuss its properties in Section 5.3.
@@
Suppose $f$ satisfies condition (2.2), and $\Theta^{\top}$ is an $f$-approximable matrix. Let $\Omega$ be the set of all the asymptotic directions for the best approximations of $\Theta^{\top}$.
@@
A more general version of Theorem 2 (without assuming (2.2)) is formulated in terms of irrationality measure functions in Theorem 2.
@@ 3.2.1 Asymptotic Dirichlet’s Theorem (due to Jarník)
As in Section 3.1, we start with a general result which follows from Jarník’s arguments. This result can be treated as analogous to the classical asymptotic corollary from Dirichlet’s theorem in the inhomogeneous case with a fixed matrix $\Theta$. It is a general version of the statement of Theorem 4 (a).
@@
Suppose $\Theta^{\top}$ is a non $f$-Dirichlet matrix. Then, for any $\boldsymbol{\eta}\in\mathbb{R}^{n}$ the pair $(\Theta,\boldsymbol{\eta})$ is $g\left(\frac{1}{\kappa^{m}}T\right)$-approximable (with $\kappa$ as in (1.3)):
@@
If we assume in addition that $f$ satisfies the condition (2.2), it implies that
@@
3.2.2 BA is null
The next theorem shows: if not all vectors $\boldsymbol{\eta}$ belong to $\widehat{\bf W}_{n,m}^{\Theta}[g]$, then the set of $g$-badly approximable shifts $\boldsymbol{\eta}$ has zero Lebesgue measure. We note that in a classical special case $g=f_{1}$ of badly approximable pairs this was shown for non-singular [^3] Non-$f_{1}$ singular in our notation $\Theta$ in [33] (Corollary 1.4).
@@ 3.2.2 BA is null
has Lebesgue measure zero in $\mathbb{R}^{n}$. If we assume that condition (2.2) holds, then the set
@@
of $g$-badly approximable shift vectors $\boldsymbol{\eta}$ has measure zero.
Note that part (b) of Theorem 4 is a reformulation of Theorem 3.
3.2.3 BA is winning
Despite being of measure zero, the set $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[g\right]$ is known to have full Hausdorff dimension, and moreover, has a stronger winning property. The dimension and winning properties of these sets were widely studied before. It was shown by Kleinbock in [35] that the set $\widehat{{\bf BA}}_{n,m}[f_{1}]$ of badly approximable pairs has full Hausdorff dimension, and by Bugeaud, Harrap, Kristensen and Velani in [6] that the set $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[f_{1}\right]$ has full Hausdorff dimension on certain fractals. It was shown by Tseng [50] that the set $\widehat{{\bf BA}}_{1,1}^{\Theta}\left[f_{1}\right]$ is winning in Schmidt’s $(\alpha,\beta)$-game. It was then shown by Moshchevitin ([42], Theorem 2) that the set $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[g\right]$ is winning, provided that the matrix $\Theta^{\top}$ is $f$-Dirichlet and $g(T)=\frac{1}{f^{-1}\left(\frac{1}{T}\right)}$. In particular, it implies that $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[f_{1}\right]$ is winning. It was shown by Einsiedler and Tseng in [17] (Theorem 1.4) and also independently by Broderick, Fishman and Kleinbock in [8] (Corollary 1.4) that the sets $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[f_{1}\right]$ are winning on certain fractals, and strengthened by Broderick, Fishman and Simmons in [10] (Theorem 1.1) who proved that the aforementioned set is HAW. We use similar techniques to prove our Theorem 4 which generalizes all these results, and proves the winning properties on certain subspaces. We note that a more general result (without the restriction (2.2)) is formulated in terms of irrationality measure functions, and can be found in Theorem 2.
@@ 3.2.3 BA is winning
Suppose $f$ satisfies condition (2.2), and $\Theta^{\top}\in{\bf D}_{m,n}\left[f\right]$ is an $f$-Dirichlet matrix. Let $\Omega$ be the set of all the asymptotic directions for the best approximations of $\Theta^{\top}$.
@@
Let us notice that winning properties of the sets $\widehat{{\bf BA}}_{n,m}^{\Theta}\,\left(=\widehat{{\bf BA}}_{n,m}^{\Theta}\left[f_{1}\right]\right)$ on affine subspaces were studied for the case of badly approximable $\Theta$ in [2]: they proved (Theorem 2.4), in the weighted setting, that the aforementioned sets are winning in the classical Schmidt’s $(\alpha,\beta)$-game on any affine subspace $\mathcal{A}$. One can ask if the winning property on arbitrary affine subspaces holds for the general $\Theta$ (and thus, if our restrictions on $\mathcal{A}$ in Theorem 4 can be omitted). The negative answer to this question is given in [16]; namely, they prove the following statement (Theorem 3.1):
@@ 3.2.4 Kurzweil-type theorems
The necessary and sufficient condition (which is also a convergence-divergence condition for a certain series) is only known when $\Theta\in\mathbb{R}$; see [20]. Theorem 3 provides a condition under which a set $\widehat{\bf W}_{n,m}^{\Theta}[g]$ has full Lebesgue measure. In this section we want to show a (non-optimal) condition under which $m\left(\widehat{\bf W}_{n,m}^{\Theta}[g]\right)=0$.
@@
$\sum\limits_{\nu=1}^{\infty}\lambda^{\frac{1}{m}}\left(Y_{\nu}^{n}\right)<\infty,$ (3.1)
@@
The function $\frac{\lambda(T)}{f(T)}$ monotonically increases. (3.2)
@@
$\tilde{g}(T)=\tilde{g}_{n,m}(T)=T^{\frac{n}{m}}\cdot\frac{f^{\frac{n}{m}}\left(H(T)\right)}{H(T)}=\frac{\lambda^{\frac{n}{m}}\left(H(T)\right)}{H(T)}.$ (3.3)
@@
Theorem 6 follows from Theorem 3; see Section 7.1 for details.
@@
Suppose that $\Theta^{\top}$ is an $f$-Dirichlet matrix, $\lambda$ is a decreasing function on $\mathbb{R}_{+}$, and conditions (2.2), (2.4), (2.5) and (2.6) are satisfied.
@@
$\sum\limits_{\nu=1}^{\infty}\left(\frac{u(Y_{\nu}^{n})}{g(Y_{\nu}^{n})}\right)^{\frac{\gamma}{m+\gamma n}}=\sum\limits_{\nu=1}^{\infty}\lambda^{\frac{1}{m}}\left(Y_{\nu}^{n}\right)<\infty,$ (3.4)
@@
Let us note that we can always find such a $u$ that satisfies conditions (2.4) and (3.4): one can always take
@@
for $\varepsilon>0$. The convergence in (3.4) is guaranteed by Proposition 1 (a).
@@ 3.3 Khintchine’s results revisited
For any pair $(\Theta,\boldsymbol{\eta})$ that is not trivially singular one has $\omega(\Theta)\geqslant\hat{\omega}(\Theta,\boldsymbol{\eta})$, meanwhile for any $\Theta$ there exists $\boldsymbol{\eta}$ with $\hat{\omega}(\Theta,\boldsymbol{\eta})=\omega(\Theta)$. In other words
@@
In Section 9 we formulate and prove a general result in terms of irrationality measure functions (Theorem 1) which leads to Theorem 8. Here we would like to make some comments on Theorem 8.
@@
It is natural to ask whether it is possible to generalize the statement (b) of Theorem 8 to affine subspaces of $\mathbb{R}^{n}$. Here we want to explain why this is not always possible in a simple low-dimensional case. One can easily construct straightforward higher dimensional generalizations of this statement.
@@
Indeed: let us show that we can find such an $\mathcal{A}$ of the form $\mathcal{A}_{\eta_{1}}=\{(\eta_{1},\eta):\,\,\eta\in\mathbb{R}\}$, where we can take $\eta_{1}$ from a HAW subset of $\mathbb{R}$.
@@
is HAW. Fix such an $\eta_{1}$; it is easy to see that, just by a simple observation that
@@ 4.1 An example: powers and logarithms
$f_{\bf a}=f_{(a,b)}=T^{-a}\left(\log T\right)^{-b}.$ (4.1)
Then, according to (1.2), the inverse function of $f_{\bf a}$ is equal to $g(T)=(1+o(1))\cdot g_{\bf a}(T)$, where
$g_{\bf a}(T)=a^{-\frac{b}{a}}T^{-\frac{1}{a}}\left(\log T\right)^{\frac{b}{a}}.$ (4.2)
Let us also notice that due to Proposition 1 (a), for any $\delta>0$ one can take $\lambda(T)=\left(\log T\right)^{-m(1+\delta)}$ in Theorem 6, and in this case the function $\tilde{g}$ in Theorem 6 satisfies
$\tilde{g}(T)\asymp\tilde{g}_{{\bf a},\delta}(T)=T^{-\frac{1}{a}}\left(\log T\right)^{\frac{b}{a}-\frac{(m+n)(1+\delta)}{a}}=\left(\log T\right)^{-\frac{(m+n)(1+\delta)}{a}}\cdot g_{\bf a}(T).$ (4.3)
@@
In a similar vein, we show some results for asymptotic inhomogeneous approximation:
@@
4.2 Criteria of bad and very good approximability
In a recent paper [19], the following two criteria for bad approximability and very good approximability were presented: